OpenStudy (shaleiah):

A quadratic function is given as f (x) = 4x^2 + 8x + 10. Write an equivalent function

OpenStudy (shaleiah):

@jojokiw3

OpenStudy (shaleiah):

@Ms-Brains

OpenStudy (anthonyym):

You can factor the equation and take a common factor out of the three terms.

OpenStudy (shaleiah):

How could I solve this by using the complete the square formula?

OpenStudy (anthonyym):

You want to solve for x? It's not an equation so I don't think you can complete the square. But I may be wrong because I kind of forget how to complete the square

OpenStudy (jojokiw3):

We have the quadratic equation \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] and we have the form equation \[ax^2 + bx + c \] How will you implement this?

OpenStudy (jojokiw3):

\[4x^2 + 8x + 10\] What is your a, b, and c?

OpenStudy (shaleiah):

a: 4 b:8 c:10

OpenStudy (jojokiw3):

Ok. Now fill in the values with the quadratic formula

OpenStudy (shaleiah):

\[-8\pm \sqrt{8^2-4(4)(10)}\] -------- 2(4)

OpenStudy (jojokiw3):

Solve it. :P Note: You'll get an imaginary number.

OpenStudy (shaleiah):

\[\frac{ 8\pm \sqrt{96} }{ 8 }\]

OpenStudy (jojokiw3):

Can you simplify further?

OpenStudy (jojokiw3):

Simplify the square root first.

OpenStudy (shaleiah):

sqrt 96= 9.7979

OpenStudy (jojokiw3):

Careful actually. You have a couple sign mistakes in your \[\frac{ 8 \pm \sqrt{96} }{ 8 }\] Find them

OpenStudy (shaleiah):

\[\frac{ 8\pm \sqrt{96i} }{ 8 }\]

OpenStudy (jojokiw3):

Okay uh. \[\frac{ -8 \pm \sqrt{64-160} }{ 8 } \neq \frac{ 8 \pm \sqrt{96} }{ 8 }\] What about the negative sign in front of the 8 and the smaller number subtracting the larger? When a small number subtracts a larger number, you get a negative.

OpenStudy (shaleiah):

You're right, sorry.

OpenStudy (jojokiw3):

So now we have \[\frac{ -8 \pm \sqrt{96} }{ 8 }\] We can reduce our square root more \[\sqrt{-96} = \sqrt{16*-6} = 4\sqrt{-6} = 4i \sqrt{6}\] Recall that \[i = \sqrt{-1}\] Now we end up with \[\frac{ -8 \pm 4i \sqrt{6} }{ 8 } \] Now you can simply do some simple cancelling, between the integers and you have your answer.

OpenStudy (jojokiw3):

lol error in my first equation...forgot the negative in the square root.

OpenStudy (jojokiw3):

\[-1 \pm 0.5i \sqrt{6} = answer\]